OpenStudy (anonymous):

Factor each trinomial below. Please show your work and check your answer. (1 point each) x2 – 8x + 15 a2 – a – 20 a2 – 5a – 20 a2 + 12ab + 27b2 2a2 + 30a + 100

OpenStudy (precal):

do you know how to factor?

OpenStudy (agent47):

Hint: factors of b that add up to c

OpenStudy (precal):

x^2-8x+15 look at the last number it is 15 can you name the factors or 15 1 times 15 anymore?

OpenStudy (agent47):

3 and 5... but 3+5 is not 8.. so what other factors do you need?

OpenStudy (anonymous):

Yes, 3+5 DOES equal 8 :)

OpenStudy (precal):

Ok let us know when you would like to learn the process, meanwhile I leave this to someone else to finish

OpenStudy (anonymous):

okay sorry i was doing something

OpenStudy (precal):

that's ok, someone else will come help

OpenStudy (anonymous):

Ok, so when you multiply two binomials (e.g. X+3), it takes the form of: \[(X+a)(X+b)\]where a,b are numbers and X is a variable. If you multiply that out, you get: \[X^{2}+bX+aX+a*b = X^{2}+(a+b)X + (a*b)\] So if you match this up with the first equation you get: \[X^{2}-8X+15\] -8 = a+b 15=a*b Possible factors of 15: 1*15=15 -1*-15=15 3*5=15 -3*-5=15 Since -3*-5=15, and -3+(-5)=-8, you know that they are a and b. Place them back into the original factor, solve for X, and you're done! \[(X+a)(X+b)\]

OpenStudy (precal):

now you can do all of the rest, good job Wired explaining it

OpenStudy (anonymous):

TYVM :) Almost went with (aX+b)(cX+d), but that's more complicated. Best to start with the basics :)

OpenStudy (precal):

yes always start with the basics